OpenStudy (anonymous):

Find the HCF of 2x^4-5x^3-2x+1 and 4x^3+8x^2+5x+3

OpenStudy (lgbasallote):

is hcf highest common factor?

OpenStudy (anonymous):

@lgbasallote yes

OpenStudy (perl):

gcf is the conventional name for it

OpenStudy (perl):

so you need to find their roots, and then look at the factored forms

OpenStudy (perl):

2x^4-5x^3-2x+1 = (x^2 -3x + 1) (2x^2 +x+1) 4x^3+8x^2+5x+3 = (2x+3)( 2x^2 + x + 1) so the gcf is (2x^2 + x + 1)

OpenStudy (perl):

but how in the world did i factor that??

OpenStudy (perl):

this is actually a tough problem, unless i am missing a trick

OpenStudy (perl):

i typed it into my TI 89, that is cheap

OpenStudy (perl):

i dont deserve a medal, somebody else do a solution without calculator

OpenStudy (perl):

well the second quartic has a rational root of -3/2

OpenStudy (perl):

you can use long division to find the cubic , then try dividing that into the other polynomial . since it is irreducible